# Fine-grained Attention I/O Complexity: Comprehensive Analysis for Backward Passes

### v1

@@ 6 Conclusion
This work provided a comprehensive analysis of the I/O complexity for attention mechanisms, focusing on backward passes. We established tight bounds on I/O complexity for both small and large caches. Our results confirm that FlashAttention is optimal for both forward and backward on large cache sizes. For small cache sizes, we provided improved upper and lower bounds compared to existing methods. Additionally, we derived lower bounds for sparse attention for both forward and backward and across cache sizes. Our findings complete the theoretical foundation for I/O complexity in attention mechanisms, offering insights for efficient LLM training and inference. We leave exploring practical implementations leveraging these theoretical insights and investigating I/O complexity for other emerging attention variants as our future work.
Acknowledgement
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
@@ Appendix A More Related Work
The exceptional success of generative large language models (LLMs), such as GPT-4 [2], Claude 3 [7], Gemini 1.5 [89], Llama 3.1 [73], Mistral Nemo [57], Phi 3.5 [4], is fundamentally attributed to the transformer architecture introduced by VSP+ [108] and all support at least 128k input token length. The transformer architecture and its self-attention mechanism have become indispensable in leading natural language processing (NLP) models [22], demonstrating remarkable capabilities across a diverse array of applications, including language translation [54], sentiment analysis [106], language modeling [76], the integration of differential privacy [90, 70], and multi-modal tasks [124, 71, 111]. Transformers’ emergent compositional abilities [30, 117] and proficiency in in-context learning [82, 74, 100] have led some to consider them as early indicators of Artificial General Intelligence (AGI) [14]. As such, the transformer architecture continues to play a pivotal role in advancing the field of AI.
@@ B.1 Notations
For any positive integer $n$, we define $[n]:=\{1,2,\dots,n\}$. For two same length vector $x$ and $y$, we use $\langle x,y\rangle$ to denote the inner product between $x$ and $y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $\circ$ to denote the Hadamard product i.e. the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$. We use $x\circ y$ to denote vector that $i$-th entry is $x_{i}y_{i}$. Let ${\bf 1}_{n}$ denote the length-$n$ all ones vector. It is not hard to see that $\langle x\circ y,{\bf 1}_{n}\rangle=\langle x,y\rangle$. For a vector $x$, we use $x^{\top}$ to denote the transpose of $x$. For a matrix $A$, we use $A^{\top}$ to denote the transpose of matrix $A$. For a matrix $A$, we use $\exp(A)$ to denote the matrix that $(i,j)$-th coordinate is $\exp(A_{i,j})$.
Given a matrix $A\in\mathbb{R}^{n\times m}$, we index an individual entry as $A[i,j]$. The $i$-th row is denoted $A[i]$ while the $j$-th column is denoted $A[*,j]$. $A[i_{1}:i_{2},j_{1}:j_{2}]$ denotes a block of $A$ consisting of entries $(i,j)$ where $i\in[i_{1},i_{2}]$ and $j\in[j_{1},j_{2}]$. Given a block size $B$, the block $A[(i-1)\cdot B+1:i\cdot B,(j-1)\cdot B+1:j\cdot B]$ is denoted $A^{(B)}[i,j]$.
@@ C.1 Algorithm for Attention Backward Without Cache
From Lemma C.1, we can prove the time complexity. Since the stored matrices have three sizes, namely $n\times d$, $n\times n$, $d\times d$, the space complexity is $O(n^{2}+nd+d^{2})=O(n^{2}+d^{2})$. ∎
@@ C.2 Algorithms for Attention Backward in Small Cache
36:     return $f$triangleright $f\in\mathbb{R}^{n\times n}$, where $f$ is defined in Definition B.5
@@
26:     return $q$triangleright $q\in\mathbb{R}^{n\times n}$, where $q$ is defined in Definiton B.8
@@
Algorithm 4 Attention gradient computation with cache phase 3. Compute $p$.
@@
8:              Compute $v^{(B)}[i]\leftarrow v^{(B)}[i]+(f^{(B)}[i,j]\circ q^{(B)}[i,j])\cdot{\bf 1}$triangleright $v=(f\circ q)\cdot{\bf 1}$
@@
20:     return $p$triangleright $p\in\mathbb{R}^{n\times n}$, where $p$ is defined in Definiton B.9
@@
3:     /* Phase 4: Compute $\frac{\mathrm{d}L(X)}{\mathrm{d}X}$/
@@
26:     return $g$triangleright $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ C.3 Upper Bound for Attention Backward in Small Cache $M=o(d^{2})$
2:     $f\leftarrow\textsc{AttentionGradientCachePhase1}(A_{1},A_{2},X,M)$triangleright see Algorithm 2
3:     $q\leftarrow\textsc{AttentionGradientCachePhase2}(A_{3},\mathrm{d}O,f,Y,M)$triangleright see Algorithm 3
4:     $p\leftarrow\textsc{AttentionGradientCachePhase3}(q,f,M)$triangleright see Algorithm 4
5:     $g\leftarrow\textsc{AttentionGradientCachePhase4}(A_{1},A_{2},p,M)$triangleright see Algorithm 5
6:     return $g$triangleright $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ D.1 Algorithms for Attention Backward in Large Cache
18:     return $S,h$triangleright $S,h\in\mathbb{R}^{n\times d}$
@@
21:              Compute $g\leftarrow g+T_{*,j}A_{2,j}$triangleright $g=TA_{2}$
@@
27:     return $g$triangleright $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ D.2 Upper Bound for Attention Backward in Large Cache $M=\Omega(d^{2})$
2:     $S,h\leftarrow\textsc{AttentionGradientLargeCachePhase1}(A_{1},A_{3},X,Y,M)$triangleright see Algorithm 7
3:     $g\leftarrow\textsc{AttentionGradientLargeCachePhase4}(A_{1},A_{2},h,S,O,\mathrm{d}O,l,M)$triangleright see Algorithm 8
4:     return $g$triangleright $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ E.1 Basic Definition in Graph Theory
[…]
@@ E.2 Previous Tools for I/O Complexity
Suppose that $M=\Omega(d^{2})$ and $A\in\mathbb{R}^{n_{1}\times d},B\in\mathbb{R}^{d\times n_{2}}$. Let $\mathcal{P}$ be an $M$-partition of the computational graph of any algorithm that computes $AB$ using standard matrix multiplication. Then for each $V^{\prime}\in\mathcal{P}$, $V^{\prime}$ contains at most $O(\frac{M^{2}}{d})$ product nodes $A_{i,k}B_{k,j}$, sum nodes $(AB)_{i,j}$, and all intermediate nodes in the summation trees.

### v2

@@ 6 Conclusion
In this work, we established tight bounds on I/O complexity for both small and large caches. Our results confirm that FlashAttention is optimal for both forward and backward on large cache sizes. For small cache sizes, we provided improved upper and lower bounds compared to existing methods. Additionally, we derived lower bounds for sparse attention for both forward and backward and across cache sizes. Our findings complete the theoretical foundation for I/O complexity in attention mechanisms and provide a deeper understanding of memory efficiency in attention computations, offering the insights for optimizing implementations in future deep learning architecture, and speeding up training and inference of large language models.
@@ Appendix A More Related Work
The exceptional success of generative large language models (LLMs), such as GPT-4 [2], Claude 3 [7], Gemini 1.5 [90], Llama 3.1 [74], Mistral Nemo [58], Phi 3.5 [4], is fundamentally attributed to the transformer architecture introduced by VSP+ [110] and all support at least 128k input token length. The transformer architecture and its self-attention mechanism have become indispensable in leading natural language processing (NLP) models [23], demonstrating remarkable capabilities across a diverse array of applications, including language translation [55], sentiment analysis [108], language modeling [77], the integration of differential privacy [91, 71], and multi-modal tasks [126, 72, 113]. Transformers’ emergent compositional abilities [31, 119] and proficiency in in-context learning [83, 75, 102] have led some to consider them as early indicators of Artificial General Intelligence (AGI) [14]. As such, the transformer architecture plays a pivotal role in advancing the field of AI.
Attention Computation Acceleration. The quadratic time complexity of attention computation with respect to the length of the input sequence [110] poses significant computational challenges, especially for long sequences. Consequently, accelerating attention computation has become a crucial research area. From a theoretical standpoint, numerous works focus on approximating the attention matrix to accelerate computation [49, 8, 9, 69, 10, 72]. Experimental approaches involve modifying model architectures and optimizing implementations to accelerate inference. Methods such as Mamba [40, 28], Linearizing Transformers [122, 81], Hopfield Models [56, 112, 52, 118, 111, 47, 48] and PolySketchFormer [125, 62] aim to improve model performance and inference speed. System-level optimizations, such as FlashAttention [27, 26, 92] and block-wise parallel decoding [97], address bottlenecks in attention mechanisms and enhance inference speed through efficient implementation strategies. Collectively, these advancements contribute to making attention mechanisms more scalable and efficient, facilitating the deployment of large-scale language models. [96] accelerates inference by compressing the input text.
@@
Sparse Attention.
Over the past few years, there has been extensive research on sparse Transformer/Attention models with weights pruning and inputs pruning, aimed at accelerating computation and training [121, 95, 16, 106, 44, 100, 93, 67, 35, 19]. In practice, the attention matrix is sparse, significantly reducing computational costs. Theoretical studies, such as YCB+ [120], have demonstrated that sparse transformers are expressive enough and can achieve universal approximation properties.
@@ B.1 Notations
For any positive integer $n$, we define $n]:=\{1,2,\dots,n\}$. For two same length vector $x$ and $y$, we use $\langle x,y\rangle$ to denote the inner product between $x$ and $y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $\circ$ to denote the Hadamard product i.e. the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$. We use $x\circ y$ to denote vector that $i$-th entry is $x_{i}y_{i}$. Let ${\bf 1}_{n}$ denote the length-$n$ all ones vector. It is not hard to see that $\langle x\circ y,{\bf 1}_{n}\rangle=\langle x,y\rangle$. For a vector $x$, we use $x^{\top}$ to denote the transpose of $x$. For a matrix $A$, we use $A^{\top}$ to denote the transpose of matrix $A$. For a matrix $A$, we use $\exp(A)$ to denote the matrix that $(i,j)$-th coordinate is $\exp(A_{i,j})$.
Given a matrix $A\in\mathbb{R}^{n\times m}$, we index an individual entry as $A[i,j]$. The $i$-th row is denoted $A[i]$ while the $j$-th column is denoted A[,j]A[*,j]. $A[i_{1}:i_{2},j_{1}:j_{2}]$ denotes a block of $A$ consisting of entries $(i,j)$ where $i\in[i_{1},i_{2}]$ and $j\in[j_{1},j_{2}]$. Given a block size $B$, the block $A[(i-1)\cdot B+1:i\cdot B,(j-1)\cdot B+1:j\cdot B]$ is denoted $A^{(B)}[i,j]$.
@@ C.1 Algorithm for Attention Backward Without Cache
From Lemma C.1, we can prove the time complexity. Since the stored matrices have three sizes, namely $\times d, n×n$$\times n$, $d\times d$, the space complexity is $O(n^{2}+nd+d^{2})=O(n^{2}+d^{2})$. ∎
@@ C.2 Algorithms for Attention Backward in Small Cache
36:  return $f$ $\triangleright$ $f\in\mathbb{R}^{n\times n}$, where $f$ is defined in Definition B.5
@@
26:  return $q$ $\triangleright$ $q\in\mathbb{R}^{n\times n}$, where $q$ is defined in Definiton B.8
@@
Algorithm 4 Attention gradient computation with cache phase 3. Compute pp.
@@
8:     Compute $v^{(B)}[i]\leftarrow v^{(B)}[i]+(f^{(B)}[i,j]\circ q^{(B)}[i,j])\cdot{\bf 1}$ $\triangleright$ $v=(f\circ q)\cdot{\bf 1}$
@@
20:  return $p$ $\triangleright$ $p\in\mathbb{R}^{n\times n}$, where $p$ is defined in Definiton B.9
@@
3:  /* Phase 4: Compute $\frac{\mathrm{d}L(X)}{\mathrm{d}X}$ */
@@
26:  return $g$ $\triangleright$ $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ C.3 Upper Bound for Attention Backward in Small Cache $M=o(d^{2})$
2:  $f\leftarrow\textsc{AttentionGradientCachePhase1}(A_{1},A_{2},X,M)$ $\triangleright$ see Algorithm 2
3:  $q\leftarrow\textsc{AttentionGradientCachePhase2}(A_{3},\mathrm{d}O,f,Y,M)$ $\triangleright$ see Algorithm 3
4:  $p\leftarrow\textsc{AttentionGradientCachePhase3}(q,f,M)$ $\triangleright$ see Algorithm 4
5:  $g\leftarrow\textsc{AttentionGradientCachePhase4}(A_{1},A_{2},p,M)$ $\triangleright$ see Algorithm 5
6:  return $g$ $\triangleright$ $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ D.1 Algorithms for Attention Backward in Large Cache
18:  return $S,h$ $\triangleright$ $S,h\in\mathbb{R}^{n\times d}$
@@
21:     Compute $g\leftarrow g+T_{*,j}A_{2,j}$ $\triangleright$ $g=TA_{2}$
@@
27:  return $g$ $\triangleright$ $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ D.2 Upper Bound for Attention Backward in Large Cache $M=\Omega(d^{2})$
2:  $S,h\leftarrow\textsc{AttentionGradientLargeCachePhase1}(A_{1},A_{3},X,Y,M)$ $\triangleright$ see Algorithm 7
3:  $g\leftarrow\textsc{AttentionGradientLargeCachePhase4}(A_{1},A_{2},h,S,O,\mathrm{d}O,l,M)$ $\triangleright$ see Algorithm 8
4:  return $g$ $\triangleright$ $g=\frac{\mathrm{d}L(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d}$, see Definition 3.2
@@ E.1 Basic Definition in Graph Theory
Let $G=(V,E)$ be a directed acyclic graph. Let $V_{1},\ldots,V_{h}\subseteq V$ be $h$ disjoint subsets of $V$. We say that there is a cyclic dependence among $\{V_{1},\ldots,V_{h}\}$ if there exists a permutation $(i_{1},\ldots,i_{h})$ of $[h]$ such that $V_{i_{1}}$ depends on $V_{i_{h}}$, and for every $j\in\{2,\ldots,h\}$, $V_{i_{j}}$ depends on $V_{i_{j-1}}$.
@@ E.2 Previous Tools for I/O Complexity
Suppose that $M=\Omega(d^{2})$ and $A\in\mathbb{R}^{n_{1}\times d},B\in\mathbb{R}^{d\times n_{2}}$. Let $\mathcal{P}$ be an $M$-partition of the computational graph of any algorithm that computes $AB$ using standard matrix multiplication. Then for each $^{\prime}\in\mathcal{P}, VV^{\prime}$ contains at most $O(\frac{M^{2}}{d})$ product nodes $A_{i,k}B_{k,j}$, sum nodes $(AB)_{i,j}$, and all intermediate nodes in the summation trees.